use the given values to find an equation of the form f(x)=ax^2 + bx + c
3. f(0)= 6, f(2)= 18, f(4)= 34
Please explain step by step!
I would really appreciate any help!
Thank you very much!

Answer by checkley77(12844)

f(x)=ax^2 + bx + c
f(0)=6
f(0)=a*0^2+b*0+c
f(0)=c thus c=6
f(2)=18
f(2)=a*2^2+b*2+6
f(2)=4a+2b+6=18 multiply by -4 & add to the f(4) equation
------------------------
f(4)=34
f(4)=a*4^2+b*4+6
f(4)=16a+4b+6=34
-------------------------
-16a-8b-24=-72
16a+4b+6=34 add.
------------------
-4b-18=-38
-4b=-38+18
-4b=-20
b=-20/-4
b=5 ans.
4a+2*5+6=18
4a+10+6=18
4a=18-10-6
4a=2
a=2/4=1/2=.5 ans.
Proof
16*.5+4*5+6=34
8+20+6=34
34=34
f(x)=a/2+5b+6